Find three consecutive integers such that the sum of the first two integers is 24 more than the third integer.

















































































































































































Answer by actuary(112)

Let x be the first number
So x+1 is the second number and x+2 is the third number
From the statement of the problem
x+(x+1)=24+(x+2)
Collect like terms on the same side of the equation.
2x+1-x=24+2
x+1=26
Solve for x
x=25
Check the result
25+26 = 51
24+(25+2)= 24+27=51 (OK)
I hope that this helps.